Overhead Variances, Four-Variance Analysis

Oerstman, Inc., uses a standard costing system and develops its overhead rates from the current annual budget. The budget is based on an expected annual output of 126,000 units requiring 504,000 direct labor hours. (Practical capacity is 524,000 hours.) Annual budgeted overhead costs total $826,560, of which $599,760 is fixed overhead. A total of 119,400 units using 502,000 direct labor hours were produced during the year. Actual variable overhead costs for the year were $260,100, and actual fixed overhead costs were $556,250.

Required:

1. Compute the fixed overhead spending and volume variances.

Fixed Overhead Spending Variance $
Fixed Overhead Volume Variance $

2. Compute the variable overhead spending and efficiency variances. Do not round intermediate calculations

Variable Overhead Spending Variance $
Variable Overhead Efficiency Variance $
Answer #1
1. Fixed Overhead Spending & Volume Variances
Fixed Overhead Spending Variance $43,510 F
Fixed Overhead Volume Variance $31,416 U
2. Variable Overhead Spending & Efficiency Variances
Variable Overhead Spending Variance $34,200 U
Variable Overhead Efficiency Variance $10,980 U
Working .1

Fixed overhead rate = Budgeted fixed overhead / Budgeted direct labor hours

= $599,760 / 504,000 Units = $1.19 per hour
Standard hour per unit = 504,000 / 126000 = 4 hours per unit
Standard hours for actual production = 119400 * 4 = 477600 hours
Budgeted fixed overhead = $599,760
Actual fixed overhead = $556,250
Fixed overhead applied = SH * Standard rate of fixed overhead
= 477600 * $1.19 = $568,344
Fixed overhead spending variance = Budgeted fixed overhead - Actual fixed overhead
= $599,760 - $556,250 = $43,510 F
Fixed overhead volume variance = Fixed overhead applied - Budgeted fixed overhead
= $568,344 - $599,760 = $31,416 U
Working 2.
Standard rate of variable overhead = ($826,560 - $599,760) / 504000 = $0.45 per hour
Actual rate of variable overhead = $260,100 / 502,000 = $0.51812749 per hour
Variable overhead spending variance = (SR - AR) * AH
= ($0.45 - $0.51812749) * 502000 = $34,200 U
Variable overhead efficiency variance = (SH - AH) * SR
= (477600 - 502000) * $0.45 = $10,980 U